What is the percent concentration of a solution made using `125 g of NH4Cl in 750. ml of water, if the density of water is 1.00 g/mL?
1
Expert's answer
2017-03-13T15:59:06-0400
m (H2O) = 750 ml x 1.00 g/mL = 750 g
m = 750 g + 125 g = 875 g
125 g / 875 g = 0.143 x 100 % = 14.3 %
Answer: 14.3 %